1/Bronco Electronics' current assets consist of cash, marketable securities, accounts receivable, and inventories. The following data were abstracted from a recent financial statement:

Inventories $ 155,000
Total assets $ 1,430,000
Current ratio 3.3
Acid-test ratio 2.30
Debt to equity ratio 1.6


Required:
Compute the long-term assets for Bronco:

Long-term assets

2/ Bronco Electronics' current assets consist of cash, marketable securities, accounts receivable, and inventories. The following data were abstracted from a recent financial statement:

Inventories $ 170,000
Total assets $ 1,520,000
Current ratio 3.7
Acid-test ratio 2.45
Debt to equity ratio 1.5


Required:
Compute the long-term liabilities for Bronco:

Long-term liabilities

3/ Spartan Sportswear's current assets consist of cash, marketable securities, accounts receivable, and inventories. The following data were abstracted from a recent financial statement:

Inventories $ 170,000
Total assets $ 845,000
Current ratio 3.40
Acid-test ratio 1.70
Debt to equity ratio 1.60


Required:
Compute the long-term liabilities for Spartan:

Long-term liabilities
Answer #1
Answer
1 Long term assets for Bronco 965000
2 Long term liabilities for Bronco 776000
3 Long term liabilities for Spartan 350000
1 long term assets = Total assets - Current assets
computation of current asset
current ratio is 3.3 (given)
current assets/current liabilities = 3.3
therefore, current assets = 3.3 current liabilities
acid test ratio is 2.3 (given)
quick assets/current liabilties = 2.3
therefore, quick assets = 2.3 current liabilities
quick assets = current assets - inventory
2.3 current liabilities= 3.3 current liabilities - 155000
(3.3 - 2.3) current liabilities = 155000
current liabilities = 155000
therefore current asset = 3.3 current liabilities
that is, 3.3 * 155000 = 465000
therefore long term asset = 1430000-465000 = 965000
2 long term liabilities = Total debt - Current liability
a) computation of total debt
total debt + total equity = total assets = 152000(given)
debt equity ratio is 1.5 (given)
debt/equity = 1.5
therefore, total debt = 1.5 equity
1.5 equity + 1 equity = 1520000
2.5 equity = 1520000
equity = 1520000/2.5 = 608000
therefore, total debt = 1.5 equity
that is 1.5 * 608000 = 912000
b) computation of current liability
current ratio is 3.7 (given)
current assets/current liabilities = 3.7
therefore, current assets = 3.7 current liabilities
acid test ratio is 2.45 (given)
quick assets/current liabilties = 2.45
therefore, quick assets = 2.45 current liabilities
quick assets = current assets - inventory
2.45 current liabilities= 3.7 current liabilities - 170000
(3.7 - 2.45) current liabilities = 170000
1.25 current liabilities = 170000
therefore, current liabilities = 170000/1.25 = 136000
therefore long term liabilties = 912000-136000 = 776000
3 long term liabilities = Total debt - Current liability
a) computation of total debt
total debt + total equity = total assets = 845000(given)
debt equity ratio is 1.6 (given)
debt/equity = 1.6
therefore, total debt = 1.6 equity
1.6 equity + 1 equity = 1520000 845000
2.6 equity = 845000
equity = 845000/2.6 = 325000
therefore, total debt = 1.6 equity
that is 1.6 * 325000 = 520000
b) computation of current liability
current ratio is 3.4 (given)
current assets/current liabilities = 3.4
therefore, current assets = 3.4 current liabilities
acid test ratio is 1.7 (given)
quick assets/current liabilties = 1.7
therefore, quick assets = 1.7 current liabilities
quick assets = current assets - inventory
1.7 current liabilities= 3.4 current liabilities - 170000
(3.4 - 1.7) current liabilities = 170000
1.7 current liabilities = 170000
therefore, current liabilities = 170000/1.7 = 100000
therefore long term liabilties = 520000-170000 = 350000